Tanya and Jason went to the library, which was 5 miles away. Tanya left on her bike at 3:00 pm. She traveled at a rate of 15 mph. Jason left at 3:15 pm by car. He traveled at 30 mph. Who got to the library first?
Answer by rfer(16322)

5/15=0.3333 hrs = 20 min
5/30=0.1666 hrs = 10 min
3:00+20=3:20 pm
3.15+10=3:25 pm